OpenStudy (anonymous):

any one solve ? 4/5 x-3.3/4x=2.1/2

ganeshie8 (ganeshie8):

\(\huge \frac{4}{5}x - \frac{3.3}{4}x = \frac{2.1}{2}\)

ganeshie8 (ganeshie8):

its like above ?

OpenStudy (anonymous):